# Learning a Shape-Conditioned Agent for Purely Tactile In-Hand Manipulation of Various Objects

### v1

@@ Abstract
Authors are with Technical University of Munich; School of Computation, Information and Technology; Department of Computer Engineering.
also with the DLR Institute of Robotics & Mechatronics, Germany.
@@ II-B Representing Shapes for In-Hand Manipulation
$\mathcal{S}=\mathcal{S}(x,R,\mathcal{M})$ (1)
@@ II-B1 Basis Point Sets (BPS)
BPS define a fixed set of $N_{b}$ basis points $\{p_{k}\in\mathbb{R}^{3}\}_{k=1,...,N_{b}}$, for each of which the vector $v_{k}\in\mathbb{R}^{3}$ to the closest point on the mesh surface $\text{surf}(\mathcal{M})$ is computed.
@@
$p^{*}_{k}=\operatorname*{arg\,min}_{p\in\text{surf}(\mathcal{M})}$ $||p-p_{k}||$
The BPS encoding vector $V$ is then simply the concatenation of these vectors (or its associated magnitudes).
$V=\left[v_{1},...,v_{N_{b}}\right]\in\mathbb{R}^{3N_{b}}$ (3)
Unlike point cloud data, the BPS encoding has intrinsic order and thus does not require permutation-invariant network architectures such as PointNet [18].
@@ II-B2 Basis Point Sets for dynamic objects
In this work, we use the BPS encoding to represent the object shape for the policy, as well as for the state estimator. To this end, we extend the notion of BPS to the dynamic case, where the basis point encoding is now additionally a function of the time-varying object pose $(x,R)$.
$v_{k}=b(k,x,R,\mathcal{M})=\left(\operatorname*{arg\,min}_{p\in\text{surf}(\mathcal{M},x,R)}||p-p_{k}||\right)-p_{k}$ (4)
The closest point is now calculated relative to the surface of $\mathcal{M}$ translated by $x$ and rotated by $R$. If $p_{k}$ lies inside of the transformed mesh, we set $v_{k}=0$. We also define:
@@
$=\Bigl{[}b(1,x_{t},R_{t},\mathcal{M}),...,b({N_{b}},x_{t},R_{t},\mathcal{M})\Bigr{]}$
@@ II-C Shape-Conditioned Agent
Our agent consists of a policy $\pi$ and a state estimator $f$ (cf. Section II-D). The policy outputs target joint angles $q_{d}\in\mathbb{R}^{12}$.
$q_{d}\sim\pi(R_{\Delta},\mathcal{S},z,\sigma)$ (6)
As input the policy receives the relative rotation to the goal $R_{\Delta}$, a representation of the object shape $\mathcal{S}=\mathcal{S}(x,R,\mathcal{M})$ and a measurement vector $z$. For our specific agent, $z$ consists of a stack of $j=6$ joint measurements $q$ and control errors $e=q_{d}-q$ sampled at $60$Hz for a window of $0.1$ seconds:
$z_{t}=[q_{t-j},e_{t-j},...,q_{t},e_{t}]\in\mathbb{R}^{6\times 24}$ (7)
@@ II-D Learning the State Estimator
Similarly to [4], the recurrent state is composed of estimated object position $\hat{x}$, estimated orientation $\hat{R}$, and a latent state vector $l$.
$s_{t}=[\hat{x}_{t},\hat{R}_{t},l_{t}]\in\mathbb{R}^{3}\times\mathrm{SO}(3)\times\mathbb{R}^{n}$ (8)
@@
For optimizing the learnable parameters of $f_{\varphi}$, we treat the basis point set calculation as part of the computation graph, enabling us to backpropagate through $\mathcal{B}$ recurrently. For supervised training, we propose a loss function based on the differences between the ground truth basis point vectors $v_{k}$ and the predicted ones.
$d_{v}(\hat{x},\hat{R},\mathcal{M})=\frac{1}{N_{b}}\sum_{k=1}^{N_{b}}\left|||v_{k}||-||b(p_{k},\hat{x},\hat{R},\mathcal{M})||\right|$ (9)
@@
$+\log\varphi\left(d(\hat{R}_{t},R_{t})|\sigma_{R_{t}}\right)$
@@ II-E Learning a Policy / Reinforcement Learning Task
Our reinforcement learning task is similar to the one described in [4], where we give a positive reward when the angle to the goal $\theta_{t}=d(R_{t},R_{g})$ is smaller than the previous $\theta_{t-1}$. We also penalize deviations of the object position $x_{t}$ and joint angles $q$ from their initial values $x_{0}$ and $\bar{q}_{0}$, respectively. In summary, the reward function is:

### v2

@@ Abstract
1 Learning AI for Dextrous Robots Lab (aidx-lab.org), Technical University of Munich, Germany, and DLR Institute of Robotics & Mechatronics (German Aerospace Center)
2 Machine Vision and Perception Group, Technical University of Munich
@@ II-B Representing Shapes for In-Hand Manipulation
$\mathcal{S}=\mathcal{S}(x,R,\mathcal{M}),$ (1)
@@ II-B1 Basis Point Sets (BPS)
BPS define a fixed set of $N_{b}$ basis points $\{p_{k}\in\mathbb{R}^{3}\}_{k=1,...,N_{b}}$, for each of which the vector $v_{k}\in\mathbb{R}^{3}$ to the closest point on the mesh surface $\text{surf}(\mathcal{M})$ is computed:
@@
$p^{*}_{k}=$ $\operatorname*{arg\,min}_{p\in\text{surf}(\mathcal{M})}||p-p_{k}||.$
The BPS encoding vector $V$ is then simply the concatenation of these vectors (or its associated magnitudes):
$V=\left[v_{1},...,v_{N_{b}}\right]\in\mathbb{R}^{3N_{b}}.$ (3)
Unlike point cloud data, the BPS encoding has intrinsic order and thus does not require permutation-invariant network architectures, such as PointNets [18].
@@ II-B2 Basis Point Sets for dynamic objects
In this work, we use the BPS encoding to represent the object shape for the policy, as well as for the state estimator. To this end, we extend the notion of BPS to the dynamic case, where the basis point encoding is now additionally a function of the time-varying object pose $(x,R)$:
$v_{k}=b(k,x,R,\mathcal{M})=\big{(}\operatorname*{arg\,min}_{p\in\text{surf}(\mathcal{M},x,R)}||p-p_{k}||\big{)}-p_{k},$ (4)
where the closest point is now calculated relative to the surface of $\mathcal{M}$ translated by $x$ and rotated by $R$. If $p_{k}$ lies inside of the transformed mesh, we set $v_{k}=0$. We also define
@@
$=\Bigl{[}b(1,x_{t},R_{t},\mathcal{M}),...,b({N_{b}},x_{t},R_{t},\mathcal{M})\Bigr{]}.$
@@ II-C Shape-Conditioned Agent
Our agent consists of a policy $\pi$ and a state estimator $f$ (cf. Section II-D). The policy outputs $N_{\text{dof}}=12$ target joint angles
$q_{d}\sim\pi(R_{\Delta},\mathcal{S},z,\sigma)\in\mathbb{R}^{N_{\text{dof}}}.$ (6)
[…]
$z_{t}=[q_{t-(N_{\text{s}}-1)},e_{t-(N_{\text{s}}-1)},...,q_{t},e_{t}]\in\mathbb{R}^{2N_{\text{dof}}\cdot N_{\text{s}}}.$ (7)
@@ II-D Learning the State Estimator
Similarly to [4], the recurrent state is composed of estimated object position $\hat{x}$, estimated orientation $\hat{R}$, and a latent state vector $l$:
$s_{t}=(\hat{x}_{t},\hat{R}_{t},l_{t})\in\mathbb{R}^{3}\times\mathrm{SO}(3)\times\mathbb{R}^{n}.$ (8)
@@
For optimizing the learnable parameters of $f_{\varphi}$, we treat the basis point set calculation as part of the computation graph, enabling us to backpropagate through $\mathcal{B}$ recurrently. For supervised training, we propose a loss function based on the differences between the ground truth basis point vectors $v_{k}$ and the predicted ones:
$d_{v}(\hat{x},\hat{R},\mathcal{M})=\frac{1}{N_{b}}\sum_{k=1}^{N_{b}}\left|||v_{k}||-||b(p_{k},\hat{x},\hat{R},\mathcal{M})||\right|.$ (9)
@@
$+\log\varphi\left(d(\hat{R}_{t},R_{t})|\sigma_{R_{t}}\right).$
@@ II-E Learning a Policy / Reinforcement Learning Task
Our reinforcement learning task is similar to the one described in [4], where we give a positive reward when the angle to the goal $\theta_{t}=d(R_{t},R_{g})$ is smaller than the previous $\theta_{t-1}$. We also penalize deviations of the object position $x_{t}$ and joint angles $q$ from their initial values $x_{0}$ and $\bar{q}_{0}$, respectively. In summary, the reward function is